Ask
your own question, for FREE!
Mathematics
8 Online
A die is a six-sided figure; each side shows a number of spots; the numbers of spots on the six sides are 1, 2, 3, 4, 5, and 6. Unless the problem says otherwise, you can assume that the die is fair, that is, each face has chance 1/6 of appearing regardless of what has appeared before. QQ A die is rolled three times. Find the chance that the face with 4 spots appears at least once.
Still Need Help?
Join the QuestionCove community and study together with friends!
P(at least once)=1-P(zero occurrences )
1/6 and 5/6
\[1-\left(\frac{5}{6}\right)^3\]
no.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
20 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal